# PHYSICISTS JOKE

### v1

@@ 4.6 Mathematization
Author: N. Vanserg (Pseudonym of Harvard Geology Professor H. E. McKinstry)
Original Source: The American Scientist, Vol. 46, No. 3 (1958).
In an article published several years ago, the author already hinted (with all the subtlety befitting the case) that, since most scientific truths (if one sorts them out) are relatively simple, any self-respecting scientist must, for the purpose of self-defense, try to prevent his colleagues from understanding that his own ideas are also simple. Therefore, if you manage to give your publications a sufficiently incomprehensible and uninteresting form, no one will attempt to read them, but everyone will bow before your erudition.
What is Mathematization?
A further development of the currently well-known methods of writing articles in a language only approximately resembling English can serve the subtle art of using mathematical symbols wherever possible. This art has only one drawback: what if a clever rogue is found, no worse than you skilled in these primitive tricks, who figures out the confused argumentation and discovers the hidden simplicity? Fortunately, there are many ways to nip such a vile attempt in the bud.
@@
The most ancient method is to write the wrong letters in formulas, for example $\phi$ instead of $\psi$. Even the simple placement of the symbol $\exp$ to the right of the brackets is sometimes capable of working wonders. This trick is conscious fraud, but it rarely entails punishment, since one can always shift the blame to the typesetter. Generally, the author, as a rule, does not have to work on inventing such traps, for typists and formula-inserters finely feel the needs of authors and themselves show initiative and good will in this regard. One can fully rely on them and then just not read the proofs.
Strategy of Secret Symbols
But what if, thanks to an accidental concurrence of circumstances, the formulas escaped distortion beyond recognition? After all, the reader, if he knows what each symbol means, will figure them out. This is where the front line of your defense passes: make it so that he never learns this. For example, you can print in black and white in a footnote on page 35 that $V_{1}$ is the total volume of phase 1, and on page 873 calmly introduce it into an equation. And your conscience will be absolutely clear—after all, you did say eventually what this symbol means. By secretly introducing all letters of the Latin, Greek, and Gothic alphabets, you can force the curious reader, interested in some paragraph, to read the entire book in reverse order to find out the meaning of the notations. Books that read from end to beginning no worse than from beginning to end produce the greatest impression.
@@
When reading backwards becomes a habit for the reader and he considers this method normal, confuse the tracks. Insert, for example, $\mu$ into an equality on page 66, and wait with its definition until page 86…
But the moment has come when the reader thinks he already knows all the letters. It is high time to use this fact to besiege him a little. Every schoolboy knows what $\pi$ is, and this will help you break away from the enemy again. The poor reader will long automatically multiply everything by 3.1416 before he understands that $\pi$ is osmotic pressure. If you are careful and do not blab ahead of time, this will cost him about an hour and a half. The same principle can, of course, be applied to any letter. Thus, you can on page 141 absolutely honestly write that $F$ is free energy, and if the astute reader is used to the fact that $F$ is free energy in Helmholtz’s definition, then he will spend a mass of his own free energy deciphering your equations before he understands that you meant Gibbs free energy all along, which the reader thinks is $G$. generally, $F$ is a wonderful letter, it can denote not only any free energy, but also fluorine, force, farad, as well as a function of an arbitrary number of real and complex variables, thereby significantly increasing the degree of randomization dSdS ($S$, as is known, denotes entropy and… sulfur).
@@
Asterisks and numerical indices, which denote footnotes at the bottom of the page, can also be used for military ruses. Designate, without saying a bad word, some pressure as $P^{*}$, so that the unsuspecting reader searches for a footnote, which on this page, needless to say, is not there at all. And when the seeker of truth reads that $S$ constitutes $10^{14}$ cal, he will think: "Wow, what a hell of an abyss of calories!"—and will continue to think so until he reads the page to the end, stumbles there upon footnote number 14, and says: "A-a-h…".
"Therefore"
@@
But the greatest success is achieved with the help of the following trick: from the finished manuscript, you tear out two pages of calculations, and in their place insert the word "therefore" and a colon. I guarantee that the reader will guess for a good two days where this "consequence" came from. Even better is to write "obviously" instead of "therefore," since there is no reader who would dare to ask anyone for an explanation of an obvious thing. By this, you not only confuse the reader but also instill in him an inferiority complex, and this is one of the main goals.
@@ 4.7 Instructions for Authors
Original Source: The Journal of Irreproducible Results, Vol. 12 (1963).
In the journal that I head, as a rule, articles are accepted that cannot be pushed through anywhere else. If your article was returned from another journal (with a small letter), iron it with an iron and send it to our Journal (with a capital letter).
Formatting the Manuscript
Text: The manuscript must be typed on standard sheets of parchment size $81/2\times 11$ inches. Typing should be done on no more than two sides of the sheet, starting every paragraph from a new page. Leave margins of $4\dots 41/2$ inches on both sides of the page.
The author must use clear, transparent, crystal English language, preferably using not less than two-syllable words consisting of not more than four letters. One should not use words understandable to your colleagues. For example, do not write "shortened", but write "reduced", not "changed", but "modified". Do not use sentences longer than 120 words without including in them at least one verb or participle.
In everything concerning spelling, use of capital letters, and other things, follow Webster’s dictionary. The editor will redo everything anyway. Abbr. s. b. red. to min. Do not make tables from data that can be listed in the text. Do not list in the text data from which a table can be made. Do not express ratios in mg/kg.
References: Name the author, his address, and the volume number. If possible, the year. (Instead of authors’ surnames, their nicknames can be cited if they are generally known.)
Illustrations: They can be made in various ways. Special clarity is not required. On the back of each photograph, briefly outline instructions for the editor (avoid unprintable words).
Conclusions: If possible, conclusions should be shorter than the main text and presented in a form allowing their use as abstracts for abstract collections.
Not Only Physicists Joke
5.1 Parkinson’s Law in Scientific Research
@@
Original Source: New Scientist (Adapted).
To an outside observer, it seems that people who apply to Funds and Trusts, to organizations bearing the names of Rockefeller, Guggenheim, and Ford, before doing so, should investigate the question of chasing subsidies from a scientific point of view. In the contrary case, disappointment awaits them. Knowing that savings are necessary, and expenses must be cut, they often think that it is only worth coming up with a plan, drawing up a budget, and you will be received with open arms.
Let us assume that a certain Dr. Grant has a plan: to determine the incidence of philately in Hong Kong. In his dreams, he sees himself in the office of the Vanderfeller Trust, before the Board of Directors, which includes Dr. Please, Mr. Giveall, Mr. Distributeall, and Mr. Riskall. They are delighted with his plan and are worried only about whether the doctor has enough with those one and a half million dollars he is asking for. They develop the impression that five million is a more real figure. "What dollars do you have in mind?" asks the timid Dr. Grant. "I meant, of course, American dollars," answers Dr. Please. Mr. Giveall writes out a check and wishes Dr. Grant all the best. The conversation is over. Such are dreams.
But what happens in reality? Dr. Grant finds himself in an office face to face with Dr. Stingy, Mr. Zilch (Mr. Fig-vam), and Mr. G. de Garantia. Stingy says that philately is a social evil rather than a disease, and therefore goes beyond their competence. de Garantia considers the entire plan politically dangerous, since it concerns Hong Kong. All together explain that this whole idea is inadmissible, unacceptable, immoral, and illegal. Dr. Grant is thrown out into the street, and the doorman is instructed not to let him in henceforth. The papers left by him are sent to the prosecutor as material evidence for initiating a criminal case on charges of extortion.
You ask, what is Dr. Grant’s mistake? There are people who want to spend money. And he offers a plan on how to spend it. So why not? His project, rejected so categorically, is no more useless than many others already implemented. Why was he kicked out? Simply because that is his plan. That is exactly exactly why they recognized it as brilliant. The main thing in the art of obtaining subsidies and grants is to convince the officials of the Fund that the project was proposed by them, and you are merely an obedient pawn, agreeing in advance with all their suggestions.
Let us assume now that you received the subsidy—maybe from the government, maybe from public charitable organizations, or most likely from a private donation. Now your immediate task is to spend and even overspend these funds as quickly as possible, in order to ask for more next time. Benefactors prefer expenses on the construction of a building—after all, a memorial plaque with the family name of the donor can be solemnly laid in its foundation, and hung on the wall. And what advertising can Dr. Grant provide them? A tombstone on his own grave?
@@
If you decide to erect a building, it is best to attach a pompous memorial arch to an already existing hospital between the laboratory and administrative buildings, and to build a cozy apartment for yourself somewhere nearby: it is very important to live close to the place of work. Furthermore, the inscription on the memorial plaque can be composed in such a way that it creates the impression (without asserting it directly) that the donor paid for the entire hospital. The composition of the inscription should be entrusted to a specialist in ambiguities. For a small, or rather relatively small fee, I myself would undertake to compose such an inscription.
@@
Building a grand entrance is the surest way to attract donations (and remember—a hospital can have more than one gate, but several). But all buildings have a common drawback: the number of scientists working in them increases very rapidly, they fill the building and overflow it, as a result of which the problem of space arises more acutely than before. It seems that the premises themselves shrink in size (this, of course, is an optical illusion, but the fact noted above is not subject to any doubt).
A similar law of continuous growth applies to the number of journals covering progress in a certain field of science. Why does this happen? For a long time, no one could give an intelligible answer to this question, so I am not without pleasure to report that I have managed to uncover the true causes of the multiplication of scientific journals and laws by which it occurs.
I will explain these laws with the following example. Let us suppose that the oldest and most respectable of all journals on clinical medicine (Journal No. 1) was published for many years by Professor A. This professor was an outstanding person (so outstanding that many probably guess his name, which I do not dare to name here). He died several years ago. If A was mistaken in anything (and who among us does not make mistakes!), it was only in that he refused to publish all articles with which he did not agree. In practice, this means—all articles written slightly above the student level. This lasted for several years and terribly bored Professor B, who never agreed with Professor A.
@@
If, for example, they were asked to write the word "vinaigrette", I am sure that they would write it differently. With such a striking difference in views, it is not surprising that the articles of Professor B for twenty-three years were invariably returned to the author. Upon the expiration of this term, he decided to found Journal No. 2. This publication began to come out on a more liberal basis, and at first, everything was printed in it, except for the works of those authors regarding whom it was accurately known that they are followers of Professor A. But Professor B had his own high principles. He believed that any views, including those which differ slightly from his own, deserve the right to free expression; he insisted only that they be set out consistently and scientifically. And so, one day, he had to reject the works presented by Professor C. (I must speak with caution: he is alive and well and deservedly receives a pension.) Everyone considered him an original and interesting thinker, but found that he was somewhat hasty in his conclusions and slightly careless in the presentation of results.
Having discovered that his articles are rejected by Journals No. 1 and No. 2, he became the founder and first publisher of Journal No. 3, which did not refuse the most unintelligible works on the vaguest topics. Everyone knows what journal I have in mind. But if you know, then you must notice that he also has a reputation that he values. His literary level is very high. Perhaps nothing is communicated in his reports, but the drawings prove the assertions that they must illustrate, but the grammar in this journal is above all criticism. Following his clinical advice, you can become a murderer, but the pages of this publication were never defiled by clumsy phrasing. Feeling obliged to protect the literary reputation of the journal (only that!), the editor was forced to reject the works of Professor D.
But we know all this—Professor D is not such a person who can be shut out of the printed pages. And here the reader receives Journal No. 4. But even D must draw a dividing line somewhere! He stubbornly refuses to publish the works of Professor E on the pretext that E does not know spelling. (And this, frankly speaking, is true.) Of course, some will argue that the article can be finalized in the editorial office. (And this, of course, is fair.) But Professor D can also be understood, and I would not blame him for narrow-mindedness. He simply does not want it to be rumored about Journal No. 4 that they accept everything that is typed on a typewriter on one side of the sheet at double intervals. He must maintain the prestige of the journal. On the other hand, no one will raise a hand to throw a stone at Professor E for the fact that he started publishing Journal No. 5. It is precisely such a development of events that led to the fact that only on questions of dentistry and prosthetics, we publish about eighty journals.
If progress in science were measured only by the quantity of published works, then the number of existing journals could become a source of satisfaction and pride. But it is necessary to remember that every journal needs an editorial board and an editorial staff, several editors with assistants, numerous reviewers, consultants, and critics. Due to the man-hours spent on academic journalism, a mass of time is lost, intended for scientific work. If everyone involved in one way or another in a certain issue read the journals published by others (and this is the best way to avoid duplication), then it is clear that they would not have time for anything else. Interestingly, those few people whose research represents at least some value usually keep each other informed of their affairs through personal correspondence.
From what has been said, one can erroneously conclude that any person who has devoted himself to scientific work ends his career as an editor. This is incorrect. Only those who fail to take an administrative position become editors.
@@
What is the normal course of events? A person who has made a significant contribution to science is persistently offered subsidies to expand the front of research. This is exactly what happened with Dr. Hodgkins, a brilliant employee of Professor Wilkins. Can one forget his speech delivered at the meeting of the American Federation of Clinical Research in 1938! According to his theory, artists creating modern abstract paintings, as a rule, suffer from color blindness, and in some cases—dementia. With this, he created a reputation for himself, and the Dalvzylkins Foundation hurried to generously subsidize his further work. Hodgkins was asked to find out whether composers writing dance music for young people really lack a musical ear (as Professor Wilkins suspected) or whether they are simply mentally underdeveloped (the opinion to which Hodgkins himself leaned). This was a grandiose project. Sector A was intended for work with artists suffering from color blindness, and Sector B—for examining mentally deficient jazz players.
From now on, Dr. Hodgkins had to deal with the organization of the work of his staff, numbering 432 people, of whom 138 had medical or scientific qualifications, 214 had secondary and higher technical education, and 80 were hired for clerical work. Well, and Dr. Hodgkins himself lost the opportunity to engage in scientific work—obviously. But not many people understand that on this path they also lose the opportunity to manage someone else’s scientific work. They will spend all their time on problems of rational use of working premises, dealing with safety engineering, compiling vacation schedules, streamlining wages, etc., etc.
Now we can formulate the **"Parkinson’s Law for Scientific Research"**. Here it is:
"Success in scientific research causes such an increase in subsidies that the continuation of research becomes impossible."

### v2

@@ 4.6 Mathematization
Author: N. Vanserg (Pseudonym of H. E. McKinstry)
Original Source: Science, Vol. 127, No. 3306 (May 9, 1958).
In a paper published some years ago, the present author intimated (with becoming subtlety) that, since most scientific truths are relatively simple, the scientist must, in self-protection, prevent his colleagues from discovering that his ideas are simple. He must, therefore, evolve a disguising technique. The objective is to make his published work sufficiently difficult to read so that no one will try to read it, but all will respect his profundity.
The Technique of Mathematization
A device which has lately been developed to a high degree of sophistication is the use of a little mathematics. If you can present your results in a form that is sufficiently unintelligible, no one will likely discover that they are simple.
@@
The first and most elementary trick is to change the letters. If you usually use $\phi$ for the angle of dip, change it to $\psi$. Even the placement of the symbol $e$ to the right of the parenthesis has a certain psychological value. This is a conscious deception, but one that is rarely punished; one can always blame the printer. In general, however, the author need not resort to such crude devices, for the typesetter will usually cooperate in introducing enough confusion to baffle the most persistent reader.
Strategy of the Secret Symbol
But what if, by some mischance, the equations don’t get badly garbled? The mathematics is apt to be all too easy to follow, provided the reader knows what the letters stand for. Here then is your firm line of defense: at all cost prevent him from finding out. Thus you may state in fine print in a footnote on page 35 that $V^{*}$ is the total volume of phase 1, and then introduce it in an equation on page 873. If you have been careful to give no clue as to what it means, you can feel reasonably safe.
By stealthily introducing all the letters of the Roman, Greek, and Gothic alphabets, you can force the curious reader to read the book backward to find out what the symbols mean. Books that read backward as well as forward are particularly impressive.
@@
When reading backward has become a habit with the reader, and he considers this method normal, confuse the trail. Introduce, for example, $\mu$ in an equation on page 66, and do not define it until page 86.
But the moment arrives when the reader thinks he knows what all the letters stand for. This is the time to bring up your reserves. Everyone knows what $\pi$ is; this will help you to throw him off the scent. The poor reader will multiply by 3.1416 for hours before he discovers that $\pi$ is the osmotic pressure. If you are careful not to give the show away, this should cost him a good hour and a half.
The same principle can be applied to any letter. Thus you can write $F$ for free energy on page 141, and the astute reader, if he is accustomed to Helmholtz’s notation, will spend a good deal of his own free energy before he realizes that you meant Gibbs’ free energy, which he thinks is $G$. In general, $F$ is a particularly happy letter, as it can be used not only for any unspecified brand of free energy but also for fluorine, force, friction, Faradays, or a function of something or other. This tends to increase the degree of randomization dSdS ($S$, as is well known, stands for entropy and… sulphur).
@@
Asterisks and primes can also be used for camouflage. Define some pressure as $P^{*}$ and say nothing about it. The reader will search for a footnote that isn’t there. And when the seeker after truth reads that $S$ is $10^{14}$ cal, he will think: "Wow, what a hell of a lot of calories!"—and will continue to think so until he reads the page to the end, stumbles on footnote 14, and says: "Ah…".
The "Hence" Trick
@@
But the most successful device is the use of the word "hence" followed by a colon. You tear out two pages of derivation and write "hence:" followed by the conclusion. I guarantee that the reader will guess for a good two days where this "consequence" came from. Even better is to write "obviously" instead of "hence," since there is no reader who would dare to ask anyone for an explanation of an obvious thing. By this, you not only confuse the reader but also instill in him an inferiority complex, and this is one of the main goals.
@@
Who Knows Best?
"The best person to decide what research work shall be done is the man who is doing the research. The next best is the head of the department. After that you leave the field of best and turn to increasingly worse groups. The first of these is the research director, who is probably wrong more than half the time. Then comes a committee, which is wrong most of the time. Finally there is the committee of company vice presidents, which is wrong all the time."
— C. E. Mees
@@ 4.7 Instructions for Authors
Original Source: The Journal of Irreproducible Results, Vol. 12, No. 1 (1963).
The Journal accepts for publication articles which have been rejected by all other journals. If your article has been returned from another journal (lower case), iron it flat and send it to The Journal (upper case).
Preparation of Manuscript
Text: Manuscripts should be typed on standard sheets of parchment, size $81/2\times 11$ inches. Typing should be on no more than two sides of the paper. Start each paragraph on a new page. Leave margins of 4 to $41/2$ inches on both sides.
The author must use clear, concise, crystal-clear English, preferably using words of not more than two syllables and four letters. Do not use words understandable to your colleagues. For example, do not write "reduced," write "modified." Do not use sentences of more than 120 words without including at least one verb.
In all matters concerning spelling, capitalization, etc., follow Webster’s Dictionary. The editor will change everything anyway. Abbr. s. b. red. to min. Do not tabulate data that can be described in the text. Do not describe in the text data that can be tabulated. Do not express ratios in mg/kg.
References: Name of author, address, and volume number. If possible, the year. (Nicknames may be used if the author is well known.)
Illustrations: These may be prepared by any method. Clarity is not required. On the back of each photograph, write brief instructions to the editor (avoid unprintable words).
Conclusions: If possible, conclusions should be shorter than the text and presented in a form suitable for use as an abstract.
It’s Not Just Physicists Who Joke
5.1 Parkinson’s Laws in Medical Research
@@
Original Source: New Scientist, Vol. 13, No. 270 (January 25, 1962).
It appears to me, as an outside observer, that the people who apply to foundations and trusts, to institutions associated with the names of Rockefeller, Guggenheim, and Ford, should do some preliminary research in grantsmanship. Without this they are liable to disappointment. Knowing that the money must be given away and that the income will become taxable if not promptly allocated and spent, they imagine too often that the originator of a scheme for expenditure should be welcomed with open arms.
Suppose that Dr. Tapfund has a plan for measuring the incidence of philately among the Chinese minority population of Hong Kong. He pictures himself in the offices of the Vanderfeller Trust, confronted by its higher executives: Dr. Grantley, Mr. Handout, Mr. Offering, and Mr. Scatterbug. They are all delighted with his research plan but question whether a million and a half dollars will suffice. They feel that five million would be justified. "Do you mean Hong Kong dollars?" asks Tapfund, his original estimate having been in that currency. "No," says Grantley, "never heard of them. I mean dollars, U.S." Tapfund hastens to assure him that U.S. dollars will do just as well. Grantley signs a check and wishes Tapfund the best of luck. The interview is over. Such is the dream.
But what is the reality? Dr. Tapfund finds himself in the offices of the Bored Foundation, confronted by Dr. Knowsleigh, Mr. Nevershed, and Mr. Knott. Knowsleigh says that Hong Kong is colonial territory and their charter precludes them spending a cent there. Nevershed thinks that philately is more of a social evil than an actual disease, and therefore outside their terms of reference. Knott thinks the scheme politically dangerous as likely to offend Nationalist China. They agree, in chorus, that the scheme is inadmissible, unacceptable, immoral, and illegal. Tapfund is thrown into the street and the janitor instructed never to admit him again.
Where did Tapfund go wrong? There are people whose job it is to spend money. He offers them a way of spending it. Why is he thrown out? The answer is that he has proposed his own plan. That is why they unanimously rejected it. The essence of grantsmanship is to persuade the Foundation executives that it was *they* who suggested the research project and that you are a humble tool in their hands.
Suppose, now, that you have gained your grant—perhaps from the government, perhaps from a public charity, most likely from a private trust. Your immediate object is to spend this money, and spend it quickly, so as to ask for more. Benefactors generally prefer to pay for a building, because their name can be put on it. What advertisement does the donor get from Dr. Tapfund? A footnote in a medical journal? A tombstone over his grave?
@@
If you decide to build, it is best to attach a pompous memorial arch to an existing hospital, locating it between the laboratory and the administration block, and building a cozy apartment for yourself nearby. The inscription on the memorial arch can be worded so as to imply (without actually stating) that the donor paid for the *entire* hospital. I would undertake to compose such an inscription myself for a small fee.
@@
But all buildings have a common drawback: the number of scientists working in them increases very rapidly, they fill the building and overflow it, as a result of which the problem of space arises more acutely than before.
@@
A similar law of continuous growth applies to the number of journals covering progress in a certain field of science. Why does this happen? I have managed to uncover the true causes of the multiplication of scientific journals. Let us suppose that the oldest and most respectable journal (Journal No. 1) was edited for many years by Professor A. He refused to publish any papers with which he disagreed. This lasted for several years and terribly bored Professor B, who never agreed with Professor A. If, for example, they were asked to define a "unit," I am sure they would write it differently. With such a striking difference in views, it is not surprising that the articles of Professor B were invariably returned to the author. Upon the expiration of twenty-three years, he decided to found Journal No. 2. This publication began to come out on a more liberal basis, and at first, everything was printed in it, except for the works of those authors regarding whom it was known that they were followers of Professor A. But Professor B had his own high principles. He believed that any views deserve free expression; he insisted only that they be set out consistently and scientifically. And so, one day, he had to reject the works presented by Professor C. Everyone considered C an original thinker, but found that he was somewhat hasty in his conclusions. Finding his papers refused by Journals 1 and 2, he became the founder of Journal No. 3, which was open from the start to the most slipshod presentation of the vaguest ideas. But even C had to draw the line somewhere! He stubbornly refuses to publish the works of Professor D on the pretext that D does not know how to spell. Professor C does not want it rumored that Journal No. 3 accepts manuscripts typed on one side of the paper at double spacing. He must maintain the prestige of the journal. On the other hand, no one will blame Professor D for founding Journal No. 4. It is precisely such a development of events that led to the fact that on questions of dentistry alone we publish about eighty journals.
@@
What is the normal course of events? A person who has made a significant contribution to science is persistently offered subsidies to expand the front of research. This is exactly what happened with Dr. Lockstock. Can one forget his speech delivered at the meeting of the American Federation of Clinical Research in 1938? According to his theory, artists creating modern abstract paintings, as a rule, suffer from color blindness. With this, he created a reputation for himself, and the Trust hurried to generously subsidize his further work. Lockstock was asked to ascertain whether the composers of modern dance music for teenagers are all tone deaf (as Dr. Barrel had suggested) or merely subnormal (as Dr. Lockstock himself had surmised). This was a grandiose project. Division A was devoted to color blindness, and Division B to the subnormal among bandleaders. From now on, Dr. Lockstock had to deal with the organization of the work of his staff, numbering 432 people. He lost the opportunity to engage in scientific work—obviously. But not many people understand that on this path they also lose the opportunity to manage someone *else’s* scientific work. They will spend all their time on problems of space, pay, and leave rosters.
We can now formulate Parkinson’s Law for Scientific Research:
"Successful research attracts the bigger grant which makes further research impossible."
In accordance with this law, we mostly end as administrators.